OpenStudy (anonymous):

How many solutions does the following system have? y = -2/3 x + 3 and 2x + 3y = 9

OpenStudy (anonymous):

I need an answer, even with previous help I'm unable to solve it.

OpenStudy (jdoe0001):

can you solve for "y" on the 2nd equation?

OpenStudy (anonymous):

How would I?

OpenStudy (anonymous):

I'm so terrible at math, all my grades are B+'s and above except math. Its such a struggle for me.

OpenStudy (jdoe0001):

well.. do you know how to isolate a variable on one side?

OpenStudy (anonymous):

Nope ._.

OpenStudy (jdoe0001):

well... http://www.youtube.com/watch?v=ldYGiXSHa_Q <---- is how you'd solve for a variable or "isolate it" so-called, and thus you can do that on the 2nd equation once you have both equations solved for "y", check them both

OpenStudy (anonymous):

Solve the second equation for y. If you compare the the RHS of each other you will discover that they are the equation for the same line. A plot is attached.

OpenStudy (anonymous):

I'm still unable to solve it. I need an answer. :/

OpenStudy (anonymous):

To find out the number of solutions ........ solve the 2nd equation and set linear lines equal and solve again. Y = - 2/3x + 3 is good. (Its in the form of Y=Mx + b Now ... -2x + 3y = 9 needs to be in the form of Y = Mx + b 2x + 3y = 9 When you get it in Y=Mx + b you will get: 3y = -2x + 9 Divide both sides by 3 (getting Y by itself.) Y = -2/3x + 3 Set your linear points equal: y = -2/3x + 3 = -2/3x + 9 Combine like terms: -2/3x + 2/3x = 9 - 3 Solve 0 = 3 0 = 3 < Since zero CLEARLY does not equal 3, your lines don't intersect at a single point. Next case... Both lines are the SAME line over lapping each other. (That means INFINITE solutions are possible)